# Metabolic rate beyond the $3/4$ law

### v1

@@ 4 Kleiber’s law as a metabolic anchoring point
Having established the functional form of the exponent $b(n)$ along development, we can now place Kleiber’s law within this new ontogenetic framework.
@@
$B(n)=a\,M^{\,b(n)},$
we can reinterpret the meaning of Kleiber’s intercept $a$. In his classical law, $B\approx 70\cdot M^{3/4}$, the constant $70$ (in kcal/day) represents the expected basal metabolic cost for a hypothetical adult mammal of $1$ kg. In our model, this reference value is not discarded; instead, it acts as a metabolic anchoring point that sets the energetic scale, while the developmental dynamics are captured by the variation of the exponent $b(n)$.
@@
$n(M)\approx\log_{\phi}\!\left(\frac{M}{M_{0}}\right).$
@@ 5 Ontogenetic Results for the Metabolic Rate $B(n)$ in mammals
Although in previous work, already accepted for publication, we have demonstrated the predictive capacity of the model by computing the exponents $b(n)$ for specific species and comparing them with an extensive compilation of empirical data [6], the present article adopts a distinct and complementary focus. Here, our goal is to explore the theoretical implications of the model in a more discursive and didactic way, detailing step by step the derivation of the equations that culminate in the central expression $B(n)\approx 70\cdot M^{\,b(n)}$.
@@
We concentrate on analyzing the behavior of this function and the biological interpretation of its dynamics, in particular how the variation of the exponent $b(n)$ modulates the relationship between mass and metabolism throughout development. When appropriate, we make qualitative comparisons with general trends reported in the literature, illustrating how the theoretical trajectory of $B(n)$ connects with known physiological patterns, without the need for a new exhaustive statistical analysis. In this way, the text aims to consolidate the theoretical foundations of the model and expand its conceptual discussion, establishing a clear bridge between the proposed mathematical structure and the principles of developmental physiology.
@@
Table 2: Body mass intervals $M$, scaling exponents $b(n)$ (simplified and refined versions), and metabolic rates predicted by the Kleiber and ontogenetic models for different mammalian species. Metabolic rates are in kcal/day.
Species $M$ (kg) $b_{\text{simp}}(n)$ $b_{\text{ref}}(n)$ $B_{\text{Kleiber}}$ $B_{\text{simp}}(n)$ $B_{\text{ref}}(n)$
Rabbit 3.55.53.5-5.5 0.8730.8990.873-0.899 0.8380.8780.838-0.878 $1.79\times 10^{2}\;-\;2.51\times 10^{2}$ $2.09\times 10^{2}\;-\;3.24\times 10^{2}$ $2.00\times 10^{2}\;-\;3.13\times 10^{2}$
Rat 0.30.50.3-0.5 0.8720.8820.872-0.882 0.8370.8540.837-0.854 $2.84\times 10^{1}\;-\;4.16\times 10^{1}$ $2.45\times 10^{1}\;-\;3.80\times 10^{1}$ $2.56\times 10^{1}\;-\;3.87\times 10^{1}$
Mouse 0.0250.040.025-0.04 0.8290.8510.829-0.851 0.7600.8010.760-0.801 $4.40\times 10^{0}\;-\;6.26\times 10^{0}$ $3.29\times 10^{0}\;-\;4.52\times 10^{0}$ $4.24\times 10^{0}\;-\;5.31\times 10^{0}$
Cat 3.55.03.5-5.0 0.8610.8690.861-0.869 0.8190.8310.819-0.831 $1.79\times 10^{2}\;-\;2.34\times 10^{2}$ $2.06\times 10^{2}\;-\;2.83\times 10^{2}$ $1.95\times 10^{2}\;-\;2.67\times 10^{2}$
Dog (medium) 253525-35 0.8770.8910.877-0.891 0.8450.8670.845-0.867 $7.83\times 10^{2}\;-\;1.01\times 10^{3}$ $1.18\times 10^{3}\;-\;1.66\times 10^{3}$ $1.06\times 10^{3}\;-\;1.53\times 10^{3}$
Horse 450600450-600 0.7610.7910.761-0.791 0.6020.6790.602-0.679 $6.84\times 10^{3}\;-\;8.49\times 10^{3}$ $7.31\times 10^{3}\;-\;1.10\times 10^{4}$ $2.77\times 10^{3}\;-\;5.39\times 10^{3}$
Cow 600800600-800 0.8060.8390.806-0.839 0.7140.7800.714-0.780 $8.49\times 10^{3}\;-\;1.05\times 10^{4}$ $1.21\times 10^{4}\;-\;1.91\times 10^{4}$ $6.74\times 10^{3}\;-\;1.29\times 10^{4}$
Elephant 400063004000-6300 0.8630.8710.863-0.871 0.8220.8360.822-0.836 $3.52\times 10^{4}\;-\;4.95\times 10^{4}$ $8.99\times 10^{4}\;-\;1.43\times 10^{5}$ $6.40\times 10^{4}\;-\;1.05\times 10^{5}$
Blue Whale 100000150000100000-150000 0.8630.8770.863-0.877 0.8220.8450.822-0.845 $3.94\times 10^{5}\;-\;5.34\times 10^{5}$ $1.45\times 10^{6}\;-\;2.42\times 10^{6}$ $9.02\times 10^{5}\;-\;1.66\times 10^{6}$
Table 2 brings together, in a single panel, the main quantities of the model for different mammalian species. The first column indicates the species; the second shows the body mass interval $M$ considered in each case. The next two columns show the intervals of the scaling exponent obtained in our previous work, in the simplified version $b_{\text{simp}}(n)$ and in the refined version $b_{\text{ref}}(n)$. The following column displays the reference metabolic rate $B_{\text{K}}$, computed from the classical Kleiber law $B_{\text{K}}=70\,M^{3/4}$. Finally, the last two columns present the basal metabolic rate intervals predicted by the ontogenetic model, $B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$, calculated from the relation $B(n)=70\,M^{b(n)}$. In this way, the table allows one to visualize simultaneously how mass, exponent, and metabolic rate are organized into different ranges across body size, with Kleiber’s law serving as a comparative reference.
@@
Some general patterns emerge clearly. First, all exponents $b(n)$ remain in the sublinear regime, with typical values between $\sim 0.6$ and $\sim 0.9$, in agreement with the literature on metabolic scaling in mammals. Moreover, for each species, the refined version $b_{\text{ref}}(n)$ is systematically smaller than the simplified version $b_{\text{simp}}(n)$, reflecting the influence of the corrective term present in the full expression of the exponent. This difference is small for small and medium-sized species (for example, rat, rabbit, cat, and dog), but becomes progressively more relevant for large species (horse, cow, elephant, and blue whale).
@@
This behavior translates directly into the intervals of $B$. For small and intermediate mammals, the ranges of $B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$ are close to each other and also comparable, in order of magnitude, to the reference rate $B_{\text{K}}$. This indicates that, in these cases, the asymptotic approximation $b_{\text{simp}}(n)=(n-1)/n$ already captures the dependence between mass and metabolism satisfactorily, and that the ontogenetic model reproduces, in a qualitative sense, the predictions of Kleiber’s law. In contrast, for large mammals, the intervals based on $b_{\text{ref}}(n)$ are noticeably smaller than those obtained with $b_{\text{simp}}(n)$ and lie closer to $B_{\text{K}}$, revealing an attenuation of the sensitivity of $B(n)$ to mass when the logarithmic correction is taken into account. In biological terms, this means that the refined version avoids overestimating the daily energetic cost of very heavy organisms, yielding predictions more consistent with the idea that metabolic scaling becomes less steep at high body masses.
@@
In summary, Table 2 shows that the model $B(n)=70\,M^{b(n)}$ generates plausible metabolic rate intervals across several orders of magnitude in body mass and is consistently aligned with Kleiber’s law, which here plays the role of a classical baseline. At the same time, the table highlights the conceptual role of the two versions of the exponent: the simplified form offers a compact and intuitive description of the asymptotic behavior, whereas the refined form introduces an essential correction to more realistically represent the metabolism of large species, bringing its predictions closer to the rates expected from classical scaling.
Finally, we would like to make explicit our perspective on the equation $M(n)\sim M_{0}\cdot\phi^{n}$, the core of this model. We are aware that proposing that the mass growth of a complex organism follows a geometric progression based on the golden ratio is a strong and, at first glance, controversial hypothesis. Our defense of this postulate, however, is not that it faithfully describes the mass at every instant of development, but rather that it captures, in a synthetic way, the structural logic of growth organized into discrete stages, in which each step amplifies the previous one by an approximately constant factor. Within the aims of this work, we did not identify an equally simple alternative and, in a certain sense, we had no conceptual choice that would simultaneously preserve stepwise recursion, the link with the Fibonacci sequence, and the possibility of obtaining an analytical expression for $b(n)$ and, consequently, for $B(n)$. We therefore view the relation $M(n)\sim M_{0}\cdot\phi^{n}$ as a powerful heuristic tool: a deliberately radical simplification that, by isolating the principle of recursion in stages, makes visible dynamic relationships that more continuous and realistic models tend to conceal. The controversy associated with this choice is not regarded here as a flaw, but as an invitation to dialogue and to a critical exploration of the extent to which fundamental mathematical principles, such as those encoded in the Fibonacci sequence and in the golden ratio, may manifest themselves in the architecture of biological growth.
22446688101002,0002{,}0004,0004{,}000nn (developmental stage)B(n)B(n)Example of $B(n)=70\,M(n)^{b(n)}$ with $M(n)\propto\phi^{n}$$B_{\text{simp}}(n)$$B_{\text{ref}}(n)$
Figure 1: Example of the metabolic rate $B(n)$ computed from $B(n)=70\,M(n)^{b(n)}$ for $n=1,\ldots,10$, assuming $M(n)\propto\phi^{n}$ and using the simplified and refined versions of $b(n)$. The plot illustrates how the choice between the simplified and refined exponent affects the predicted metabolic trajectory, particularly in early developmental stages.
Figure 1 illustrates, in a simple way, how the metabolic rate $B(n)$ evolves along developmental stages in the context of the proposed model. For purely illustrative purposes, we assume a mass that grows as $M(n)\propto\phi^{n}$ (with $M_{0}=1$), so as to emphasize the role of the dynamic exponent $b(n)$ in the expression $B(n)=70\,M(n)^{b(n)}$. The solid curve shows the values obtained with the simplified version of the exponent, $b_{\text{simp}}(n)=(n-1)/n$, whereas the dashed curve corresponds to the refined version $b_{\text{ref}}(n)$.
Several important aspects can be observed. First, both curves describe a monotonic increase of $B(n)$ with developmental stage, reflecting the fact that more developed organisms, with larger mass, require more energy to sustain their vital functions. Second, the difference between the two versions is more evident in the early stages, when the corrective term in the refined expression still exerts a significant influence on $b(n)$. From $n\approx 4$ onward, the two trajectories become very close, indicating that, as development progresses, the asymptotic form $b_{\text{simp}}(n)$ already captures well the essential dynamics of the model. Thus, the figure serves as a visual “proof of concept”: it shows how, even in a simplified scenario, the combination of stage-structured growth with a variable scaling exponent produces a metabolic trajectory $B(n)$ that increases and is consistent with the idea of a scaling that becomes progressively less sublinear along development.
Conclusion
@@
This work started from a simple yet fundamental observation: organisms are not static, they grow and develop. Classical laws of metabolic scaling, such as Kleiber’s law and the WBE theory, were essential to describe how much the metabolism of an adult elephant differs from that of an adult rat, but they showed limitations in explaining how the metabolism of a single individual changes over the course of its own life, from the juvenile phase to maturity.
@@
To address this gap, we proposed a change in perspective. Instead of seeking a single fixed exponent that describes all organisms, we developed a model in which the scaling exponent "grows" together with the organism. Inspired by stage-like growth patterns, so common in nature and mathematically described by the Fibonacci sequence, we structured development as a succession of discrete stages. In this view, the relationship between mass and metabolism is no longer governed by an immutable rule, but is described by a dynamic power law,
$B(n)=70\,M^{b(n)},$
in which the exponent $b(n)$ varies systematically with the ontogenetic stage $n$.
@@
Ultimately, this work does not discard the achievements of previous theories, but places them in a broader and more dynamic context. By translating the recursive logic of Fibonacci into the language of physiology, we offer an alternative and complementary narrative for the phenomenon of scaling, intrinsically linked to time and to the process of becoming a complete organism. This theoretical framework, centered on principles of stepwise organization, opens the way for future investigations into how life history, energetic priorities at each phase, and the very hierarchical architecture of living beings intertwine to define the pace of life, and how these ideas can be systematically confronted with metabolic data across development.

### v2

@@ 4 Kleiber’s law as a metabolic anchoring point
Having revisited the functional form of the exponent $b(n)$ along development, we can now place Kleiber’s law within this new ontogenetic framework.
@@
$B(n)=B_{0}\,M^{b(n)},$ (11)
we can reinterpret the meaning of Kleiber’s intercept $B_{0}$. In his classical law, $B\approx 70\cdot M^{3/4}$, the constant $70$ (in kcal day-1) represents the expected basal metabolic cost for a hypothetical adult mammal of $1$ kg. In our model, this reference value is not discarded; instead, it acts as a metabolic anchoring point that sets the energetic scale, while the developmental dynamics are captured by the variation of the exponent $b(n)$.
@@
$n(M)\approx\log_{\phi}\!\left(\frac{M}{M_{0}}\right).$ (13)
@@ 5 Ontogenetic results for the metabolic rate $B(n)$ in mammals
In our previous work [8], we demonstrated the predictive capacity of the Fibonacci-based model by computing the exponents $b(n)$ for nine specific mammalian species, selected based on the availability of reliable data regarding birth and adult mass, and by comparing these exponents with empirical metabolic information from the literature. As stated in Introduction 1, our objective here is merely to explore the theoretical implications of the model in a more discursive and didactic manner, detailing step by step the derivation of the equations that culminate in the central expression $B(n)\approx 70\,M^{b(n)}$.
@@
We concentrate on analyzing the behavior of this function and the biological interpretation of its dynamics, in particular how the variation of the exponent $b(n)$ modulates the relationship between mass and metabolism throughout development. To this end, and building directly upon the stage-dependent exponents $b_{\text{simp}}(n)$ and $b_{\text{ref}}(n)$ calculated for those nine mammalian species in our previous work [8], we now compute the corresponding basal metabolic rates via $B(n)=70\,M^{b(n)}$. The results are summarized in Table 1, which also includes, for comparison, the reference metabolic rate obtained from Kleiber’s classical law, $B_{\text{K}}=70\,M^{3/4}$. When appropriate, we make qualitative comparisons with general trends reported in the literature, illustrating how the theoretical trajectory of $B(n)$ connects with known physiological patterns, without the need for a new exhaustive statistical analysis. In this way, the text aims to consolidate the theoretical foundations of the model and expand its conceptual discussion, establishing a clear bridge between the proposed mathematical structure and the principles of developmental physiology.
@@
Table 1: Body mass intervals $M$, reference metabolic rates from Kleiber’s law ($B_{\text{K}}$), and basal metabolic rate intervals predicted by the ontogenetic model ($B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$) for nine mammalian species. All metabolic rates are in kcal day-1. The underlying stage-dependent exponents $b_{\text{simp}}(n)$ and $b_{\text{ref}}(n)$ were obtained in [8].
Species Rabbit Rat Mouse Cat Dog (medium) Horse Cow
$M$ (kg) 3.53.55.55.5 0.30.30.50.5 0.0250.0250.040.04 3.53.55.05.0 25253535 450450600600 600600800800
Species/$M$ (kg)       Elephant/4000400063006300      Blue Whale /100000100000150000150000
Species $B_{\text{K}}$ $B_{\text{simp}}(n)$ $B_{\text{ref}}(n)$
Rabbit $1.79\times 10^{2}$$2.51\times 10^{2}$ $2.09\times 10^{2}$$3.24\times 10^{2}$ $2.00\times 10^{2}$$3.13\times 10^{2}$
Rat $2.84\times 10^{1}$$4.16\times 10^{1}$ $2.45\times 10^{1}$$3.80\times 10^{1}$ $2.56\times 10^{1}$$3.87\times 10^{1}$
Mouse $4.40\times 10^{0}$$6.26\times 10^{0}$ $3.29\times 10^{0}$$4.52\times 10^{0}$ $4.24\times 10^{0}$$5.31\times 10^{0}$
Cat $1.79\times 10^{2}$$2.34\times 10^{2}$ $2.06\times 10^{2}$$2.83\times 10^{2}$ $1.95\times 10^{2}$$2.67\times 10^{2}$
Dog (medium) $7.83\times 10^{2}$$01\times 10^{3} 1.1$$18\times 10^{3}$$1.66\times 10^{3}$ $1.06\times 10^{3}$$1.53\times 10^{3}$
Horse $6.84\times 10^{3}$$8.49\times 10^{3}$ $7.31\times 10^{3}$$1.10\times 10^{4}$ $2.77\times 10^{3}$$5.39\times 10^{3}$
Cow $8.49\times 10^{3}$$1.05\times 10^{4}$ $1.21\times 10^{4}$$1.91\times 10^{4}$ $6.74\times 10^{3}$$1.29\times 10^{4}$
Elephant $3.52\times 10^{4}$$4.95\times 10^{4}$ $8.99\times 10^{4}$$1.43\times 10^{5}$ $6.40\times 10^{4}$$1.05\times 10^{5}$
Blue Whale $3.94\times 10^{5}$$5.34\times 10^{5}$ $1.45\times 10^{6}$$2.42\times 10^{6}$ $9.02\times 10^{5}$$1.66\times 10^{6}$
Table 1 summarizes, in a compact form, the main quantities of the model for the nine mammalian species studied. The upper part of the table lists the species and the corresponding body mass intervals $M$ used in the calculations, while the lower part shows, for each of these species (first column), the reference metabolic rate $B_{\text{K}}$ (second column), computed from the classical Kleiber’s law and used as a comparative baseline, followed by the basal metabolic rate intervals predicted by the ontogenetic model, $B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$ (third and fourth columns, respectively), calculated from the relation $B(n)=70\,M^{b(n)}$ using the stage-dependent exponents determined in [8]. The table allows for a direct visual comparison of how the predictions of the stage-dependent scaling model relate to the classical expectation across a broad range of body sizes.
@@
As discussed in [8], for each species the refined exponent $b_{\text{ref}}(n)$ is systematically smaller than the simplified version $b_{\text{simp}}(n)$, reflecting the influence of the corrective term present in the full expression of the exponent. This difference is small for small and medium-sized species (for example, rat, rabbit, cat, and dog), but becomes progressively more relevant for large species (horse, cow, elephant, and blue whale).
@@
This behavior is directly reflected in the metabolic rate intervals shown in Table 1. For small and intermediate mammals, the ranges of $B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$ are close to each other and also comparable, in order of magnitude, to the reference rate $B_{\text{K}}$. This indicates that, in these cases, the asymptotic approximation $b_{\text{simp}}(n)=(n-1)/n$ already captures the dependence between mass and metabolism satisfactorily, and that the ontogenetic model reproduces, in a qualitative sense, the predictions of Kleiber’s law. In contrast, for large mammals, the metabolic rates based on $b_{\text{ref}}(n)$ are noticeably smaller than those obtained with $b_{\text{simp}}(n)$ and lie closer to $B_{\text{K}}$, revealing an attenuation of the sensitivity of $B(n)$ to mass when the logarithmic correction is taken into account. In biological terms, this means that the refined version avoids overestimating the daily energetic cost of very heavy organisms, yielding predictions more consistent with the idea that metabolic scaling becomes less steep at high body masses.
To illustrate the ontogenetic trajectory predicted by the model, Fig. 1 illustrates the behavior of $B(n)$ for $n=1,\dots,10$ under the idealized assumption $M(n)\propto\phi^{n}$. The two curves, corresponding to $B_{\text{simp}}(n)$ and $B_{\text{ref}}(n)$, both increase monotonically with the developmental stage, reflecting the higher energetic demands of larger, more developed organisms. The refined trajectory remains slightly below the simplified one, especially at early stages, and gradually converges toward it as $n$ increases, in agreement with the analytical behavior discussed in the previous sections.
@@
In summary, Table 1 and Fig. 1 together demonstrate that the model $B(n)=70\,M^{b(n)}$ generates plausible metabolic rate intervals across several orders of magnitude of body mass and remains consistently aligned with Kleiber’s law, which serves here as a classical baseline. At the same time, these results highlight the conceptual role of the two versions of the exponent, given that the simplified form offers a compact and intuitive description of the asymptotic behavior, whereas the refined form introduces an essential correction to more realistically represent the metabolism of large species, bringing its predictions closer to the rates expected from classical scaling.
Figure 1: Metabolic rate $B(n)$ computed from $B(n)=70\,M(n)^{b(n)}$ for $n=1,\dots,10$, assuming $M(n)\propto\phi^{n}$. Open circles and solid line denote $B_{\text{simp}}(n)$; filled squares and dashed line denote $B_{\text{ref}}(n)$
6 Conclusion
@@
This work started from a simple yet fundamental observation: organisms are not static, they grow and develop. Classical laws of metabolic scaling, such as Kleiber’s law and the WBE theory, were essential in describing how much the metabolism of an adult elephant differs from that of an adult rat, but they showed limitations in explaining how the metabolism of a single individual changes over the course of its own life, from the juvenile phase to maturity.
@@
To address this gap, we propose a change in perspective. Instead of seeking a single fixed exponent that describes all organisms, we developed a model in which the scaling exponent ”grows” together with the organism. Inspired by stage-like growth patterns, so common in nature and mathematically described by the Fibonacci sequence, we structured development as a succession of discrete stages. In this view, the relationship between mass and metabolism is no longer governed by an immutable rule, but is described by a dynamic power law $B(n)=70\,M^{b(n)}$, in which the exponent $b(n)$ varies systematically with the ontogenetic stage $n$.
@@
This work does not discard the achievements of previous theories, but places them in a broader and more dynamic context. By translating the recursive logic of Fibonacci into the language of physiology, we offer an alternative and complementary narrative for the phenomenon of scaling, intrinsically linked to time and to the process of becoming a complete organism.
We recognize that the basis of our model, the postulate that mass accumulates according to a geometric progression $M(n)\sim M_{0}\cdot\phi^{n}$ (derived from the Fibonacci recursion), is a fundamental premise, and may seem controversial to some. Within the objectives of this work, we did not identify an equally simple alternative and, in a certain sense, we did not have a conceptual choice that simultaneously preserved stepwise recursion, the connection with the Fibonacci sequence, and the possibility of obtaining an analytical expression for $b(n)$ and, consequently, for $B(n)$. We therefore view this postulate not as a literal description of biological growth, but as a deliberate heuristic that isolates the principle of stage-organized recursion, making visible dynamic relationships that more continuous models may obscure.
We believe that this theoretical framework, centered on principles of gradual organization, paves the way for future investigations into how life history, energy priorities at each stage, and the hierarchical architecture of living beings themselves intertwine to define the rhythm of life.